A source of sound S is moving with a velocity 50m/a towards stationary observer. The observer measures the frequency of the sourseas 1000hz. The apparent frequency of the source when it is moving away from the observer after crossing him is (velocity os sound =350

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
By using,
 
napproaching / nreceding = (v + vs )/ (v –vs )
 
1000 / nr = (350 + 50) / (350 – 50)
 1000/ nr = 4 / 3
 
nr = 3000/ 4 = 750 Hz
 
 
Regards
Arun (askIITians forum expert)